# Elephants Never Forget: Memorization and Learning of Tabular Data in Large Language Models

### v1

@@ Abstract
While many have shown how Large Language Models (LLMs) can be applied to a diverse set of tasks, the critical issues of data contamination and memorization are often glossed over. In this work, we address this concern for tabular data. Specifically, we introduce a variety of different techniques to assess whether a language model has seen a tabular dataset during training. This investigation reveals that LLMs have memorized many popular tabular datasets verbatim. We then compare the few-shot learning performance of LLMs on datasets that were seen during training to the performance on datasets released after training. We find that LLMs perform better on datasets seen during training, indicating that memorization leads to overfitting. At the same time, LLMs show non-trivial performance on novel datasets and are surprisingly robust to data transformations. We then investigate the in-context statistical learning abilities of LLMs. Without fine-tuning, we find them to be limited. This suggests that much of the few-shot performance on novel datasets is due to the LLM’s world knowledge. Overall, our results highlight the importance of testing whether an LLM has seen an evaluation dataset during pre-training. We make the exposure tests we developed available as the tabmemcheck Python package.
@@ 3 Testing Language Models for Memorization of Tabular Data
There are various systematic ways to test what a language model knows about a tabular dataset. For example, we can test whether the LLM can list a dataset’s feature names or tell the possible values of categorical features in the data. It is also possible to heuristically assess whether the LLM has learned different aspects of the data distribution. In this Section, we focus on tests for memorization. Additional heuristics are described in Supplement D.
@@ 3.1 Testing for memorization
We introduce four different tests to detect memorization. These tests use the fact that most tabular datasets have a canonical representation as a CSV file (which is nothing but a text document that might end up in the pre-training data). The tests extract uniquely identifying information (’canaries’) about the CSV file from the LLM (Carlini et al., 2021). Our tests are [^2] It is helpful to think about the proposed memorization tests in terms of power and significance, as we would about a statistical hypothesis test. All the tests are highly significant. This means they provide definitive evidence of memorization (though not necessarily of the entire dataset). Conversely, little can be said about the power of the tests. It seems possible that an LLM has memorized a tabular dataset, but we cannot extract it via prompting.
@@
    The Header Test: We prompt the model with the initial rows of the CSV file and ask it to complete the next rows verbatim.
@@
    The Row Completion Test: We prompt the model with a number of contiguous rows from a random position of the CSV file and ask it to complete the next row verbatim.
@@ 3.2 LLMs have memorized many of the popular tabular datasets
Table 2: Testing LLMs for memorization of tabular data. GPT-3.5 and GPT-4 have memorized many of the popular tabular datasets. The table depicts results with GPT-3.5/ GPT-4 (depicted in the table as */*) on various memorization tests. The table depicts the results of the tests in simplified form, that is, [Uncaptioned image]= evidence of memorization, [Uncaptioned image]= no evidence of memorization, [Uncaptioned image]= ambiguous result, and - = test cannot be conducted. The detailed, quantitative test results can be found in Supplement D.
@@
Table 2 shows the results of the memorization tests on 11 tabular datasets. The Table depicts the results of the tests in simplified form, where [Uncaptioned image]indicates evidence of memorization. The detailed, quantitative test results can be found in Supplement D.
@@
We see that the header test indicates memorization of all the tabular datasets that were publicly available on the internet before 2021. This means that GPT-3.5 and GPT-4 have memorized the initial rows of these datasets verbatim. On four of the publicly available datasets ( Titanic , Diabetes , Wine , Iris ), the row completion and feature completion tests also provide evidence of memorization. This means that the LLMs have not only memorized the initial rows (which might be included in papers, web pages and notebooks), but also random rows from these datasets. The same holds for the first token test. We observe that the results of the latter three tests are overall consistent, with a tendency of GPT-4 to exhibit stronger evidence for memorization than GPT-3.5. We note that this aligns with previous works that show that larger models exhibit more memorization (Carlini et al., 2022a). Interestingly, we find no evidence of memorization of random rows on Adult and Housing .
@@
There is no evidence of memorization of the novel datasets. This is to be expected for datasets released after the model was trained. In the case of the FICO dataset, it suggests that this dataset might have been protected from inclusion in the LLM training data by the need to register prior to access. [^3] Of course, it is possible that the LLM saw the dataset during training, but there was no memorization, or our memorization tests are not powerful enough.
@@ 4 Few-Shot Learning with LLMs and Tabular Data
In the previous Section, we demonstrated that GPT-3.5 and GPT-4 have memorized many popular tabular datasets. In this Section, we investigate the few-shot learning performance of GPT-3.5 and GPT-4 on the novel and memorized datasets.
@@
System: You are a classification assistant who is an expert in tabular data, data science, and cross-sectional population surveys. […] User: IF Age = 30, WorkClass = Private, fnlwgt = 196945, […] THEN Income = Assistant: $>$50k [few-shot examples] User: IF Age = 28, WorkClass = Self-emp-inc, fnlwgt = 79135, […] THEN Income = Assistant: [Model Response]
We ask the LLM to predict the value of the target variable, given the values of the other features in the data. We select 20 few-shot examples randomly and stratify the labels of the few-shot examples to match the label occurrence in the dataset. All experiments are conducted at temperature 00. A full example prompt is given in Supplement F.
@@ 4.1 Dataset Format: Original, Perturbed, Task and Statistical
Original
$0.81_{\scriptscriptstyle.01}\quad 0.96_{\scriptscriptstyle.01}$
$0.78_{\scriptscriptstyle.01}\quad 0.81_{\scriptscriptstyle.01}$
$0.74_{\scriptscriptstyle.02}\quad 0.74_{\scriptscriptstyle.02}$
$0.88_{\scriptscriptstyle.02}\quad 0.96_{\scriptscriptstyle.01}$
$0.98_{\scriptscriptstyle.01}\quad 0.99_{\scriptscriptstyle.01}$
Perturbed
$0.78_{\scriptscriptstyle.01}\quad 0.82_{\scriptscriptstyle.01}$
$0.78_{\scriptscriptstyle.01}\quad 0.81_{\scriptscriptstyle.01}$
$0.73_{\scriptscriptstyle.02}\quad 0.73_{\scriptscriptstyle.02}$
$0.88_{\scriptscriptstyle.02}\quad 0.95_{\scriptscriptstyle.02}$
$0.95_{\scriptscriptstyle.02}\quad 0.95_{\scriptscriptstyle.02}$
Task
$0.77_{\scriptscriptstyle.01}\quad 0.80_{\scriptscriptstyle.01}$
$0.75_{\scriptscriptstyle.01}\quad 0.79_{\scriptscriptstyle.01}$
$0.70_{\scriptscriptstyle.02}\quad 0.73_{\scriptscriptstyle.02}$
$0.87_{\scriptscriptstyle.03}\quad 0.87_{\scriptscriptstyle.03}$
$0.95_{\scriptscriptstyle.02}\quad 0.95_{\scriptscriptstyle.02}$
Statistical
$0.61_{\scriptscriptstyle.02}\quad 0.65_{\scriptscriptstyle.02}$
$0.70_{\scriptscriptstyle.01}\quad 0.63_{\scriptscriptstyle.02}$
$0.68_{\scriptscriptstyle.02}\quad 0.62_{\scriptscriptstyle.02}$
$0.86_{\scriptscriptstyle.03}\quad 0.90_{\scriptscriptstyle.02}$
$0.87_{\scriptscriptstyle.03}\quad 0.92_{\scriptscriptstyle.02}$
LR / GBT
0.79   /   0.80
0.86   /   0.87
0.78   /   0.75
0.98   /   0.96
0.97   /   0.95
@@
Original
$0.58_{\scriptscriptstyle.02}\quad 0.67_{\scriptscriptstyle.01}$
$0.78_{\scriptscriptstyle.01}\quad 0.78_{\scriptscriptstyle.01}$
$0.69_{\scriptscriptstyle.05}\quad 0.69_{\scriptscriptstyle.05}$
$0.58_{\scriptscriptstyle.02}\quad 0.67_{\scriptscriptstyle.01}$
$0.54_{\scriptscriptstyle.02}\quad 0.62_{\scriptscriptstyle.02}$
Perturbed
$0.57_{\scriptscriptstyle.02}\quad 0.67_{\scriptscriptstyle.01}$
$0.77_{\scriptscriptstyle.01}\quad 0.78_{\scriptscriptstyle.01}$
$0.69_{\scriptscriptstyle.05}\quad 0.71_{\scriptscriptstyle.05}$
$0.58_{\scriptscriptstyle.02}\quad 0.67_{\scriptscriptstyle.02}$
$0.54_{\scriptscriptstyle.02}\quad 0.62_{\scriptscriptstyle.02}$
Task
$0.59_{\scriptscriptstyle.02}\quad 0.65_{\scriptscriptstyle.02}$
$0.77_{\scriptscriptstyle.01}\quad 0.77_{\scriptscriptstyle.01}$
$0.67_{\scriptscriptstyle.05}\quad 0.71_{\scriptscriptstyle.05}$
$0.61_{\scriptscriptstyle.02}\quad 0.68_{\scriptscriptstyle.01}$
$0.54_{\scriptscriptstyle.02}\quad 0.65_{\scriptscriptstyle.01}$
Statistical
$0.63_{\scriptscriptstyle.02}\quad 0.66_{\scriptscriptstyle.01}$
$0.59_{\scriptscriptstyle.02}\quad 0.57_{\scriptscriptstyle.02}$
$0.56_{\scriptscriptstyle.05}\quad 0.55_{\scriptscriptstyle.05}$
$0.60_{\scriptscriptstyle.02}\quad 0.59_{\scriptscriptstyle.02}$
$0.52_{\scriptscriptstyle.02}\quad 0.48_{\scriptscriptstyle.02}$
LR / GBT
0.78   /   0.78
0.80   /   0.80
0.76   /   0.66
0.70   /   0.69
0.64   /   0.67
@@ 4.2 Memorization leads to inflated performance estimates
Table 3 depicts the few-shot learning performance of GPT-4 and GPT-3.5 on 10 different tabular datasets. On the memorized datasets depicted in Panel A of Table 3, the performance is quite impressive. In particular, GPT-4 outperforms logistic regression on 3 out of 5 datasets when prompted with the original data. However, the predictive performance of GPT-4 and GPT-3.5 drops as we move from the original data to the perturbed data. As depicted in Figures 1(a) and 1(c), the accuracy of GPT-4 and GPT-3.5 on the memorized datasets consistently drops as we present the data in the original, perturbed, task and statistical format. In particular, GPT-4 on the task data outperforms logistic regression only on 1 out of 5 datasets, and even there, the measured difference lies within the standard error.
Panel B of Table 3 depicts the few-shot learning performance of GPT-4 and GPT-3.5 on novel datasets. In contrast to the results in Panel A, GPT-4 in the original format does not outperform logistic regression on any dataset. Moreover, we also don’t observe significant changes between the original, perturbed, and task formats. This is again depicted in Figures 1(b) and 1(d). While it is always possible that the few-shot learning performance of an LLM varies with subtleties such as the length of a feature name or the number of significant digits of a number, the results in Panel B of Table 3 indicate that GPT-4 and GPT-3.5 are fairly robust to such changes.
@@
4.3 Performance on unseen datasets depends on feature names and variable format
@@
4.4 Language models have a limited ability to act as statistical predictors

### v3

@@ Abstract
While many have shown how Large Language Models (LLMs) can be applied to a diverse set of tasks, the critical issues of data contamination and memorization are often glossed over. In this work, we address this concern for tabular data. Specifically, we introduce a variety of different techniques to assess whether a language model has seen a tabular dataset during training. This investigation reveals that LLMs have memorized many popular tabular datasets verbatim. We then compare the few-shot learning performance of LLMs on datasets that were seen during training to the performance on datasets released after training. We find that LLMs perform better on datasets seen during training, indicating that memorization leads to overfitting. At the same time, LLMs show non-trivial performance on novel datasets and are surprisingly robust to data transformations. We then investigate the in-context statistical learning abilities of LLMs. While LLMs are significantly better than random at solving statistical classification problems, the sample efficiency of few-shot learning lags behind traditional statistical learning algorithms, especially as the dimension of the problem increases. This suggests that much of the observed few-shot performance on novel real-world datasets is due to the LLM’s world knowledge. Overall, our results highlight the importance of testing whether an LLM has seen an evaluation dataset during pre-training. We release the tabmemcheck Python package to test LLMs for memorization of tabular datasets.
@@ 3 Testing Language Models for Memorization of Tabular Data
There are various systematic ways to test what a language model knows about a tabular dataset. For example, we can test whether the LLM can list a dataset’s feature names or tell the possible values of categorical features in the data. It is also possible to heuristically assess whether the LLM has learned different aspects of the data distribution. In this Section, we focus on tests for verbatim memorization (Carlini et al., 2019; 2021). Additional heuristics and details are described in Supplement D.
@@ 3.1 Testing for memorization
We introduce four different tests to detect memorization. These tests use the fact that most tabular datasets have a canonical representation as a CSV file (which is nothing but a text document that might end up in the pre-training data). Intuitively, we say that a tabular dataset is memorized if the LLM can consistently generate it. In our context of tabular datasets, this is justified by the fact that these datasets contain random variables: It is impossible to consistently reproduce the realizations of random variables unless the values of the random variables have been seen before (i.e., during pre-training). [^2] This approach has been lightly formalized in Carlini et al. (2019) and Carlini et al. (2021), where the authors use the term ’canary’.
Previous work on memorization in language models has often relied on the prefix-suffix pattern, where the model is provided with the initial part of a string as a prefix and asked to provide a completion (Carlini et al., 2019; 2021; 2022b). In this work, we use few-shot learning to condition chat models on the task of completing a given prefix, a prompt strategy that works well for GPT-3.5 and GPT-4, as well as for many other LLMs. The different tests can be performed with the tabmemcheck Python package. The tests are
@@
    The Header Test: We prompt the model with the initial rows of the CSV file and ask it to complete the next rows verbatim. This test measures the memorization of the initial rows of the dataset.
@@
    The Row Completion Test: We prompt the model with a number of contiguous rows from a random position of the CSV file and ask it to complete the next row verbatim. This
@@
It is helpful to think about the proposed memorization tests in terms of power and significance, as we would about a statistical hypothesis test. All the tests are highly significant. This means they provide definitive evidence of memorization (though not necessarily of the entire dataset). Conversely, little can be said about the power of the tests. It seems possible that an LLM has memorized a tabular dataset, but we cannot extract it via prompting.
@@ 3.2 LLMs have memorized many of the popular tabular datasets
Table 2: GPT-3.5 and GPT-4 have memorized many of the popular tabular datasets. The table depicts the results of different memorization tests with GPT-3.5 and GPT-4 (depicted in the table as */*). Test results are depicted in simplified form, that is, [Uncaptioned image]= evidence of memorization, [Uncaptioned image]= no evidence of memorization, [Uncaptioned image]= ambiguous result, and - = test cannot be conducted. The detailed, quantitative test results can be found in Supplement D.
@@
Table 2 shows the results of the memorization tests on 11 tabular datasets for GPT-3.5 and GPT-4. The Table depicts the results of the tests in simplified form, where [Uncaptioned image]indicates that the model is able to generate the respective parts of the dataset. Quantitative test results and model generations can be found in Supplement D.
@@
The header test indicates memorization of all the tabular datasets that were publicly available on the internet before 2021. This means that GPT-3.5 and GPT-4 have memorized the initial rows of these datasets. On Titanic , Diabetes , Wine , and Iris , the row completion, feature completion, and first token test also provide evidence of memorization. This means that GPT-3.5 and GPT-4 have not only memorized the initial rows of these datasets (which are frequently in Jupyter notebooks) but also random rows. Of course, this means that GPT-3.5 and GPT-4 have essentially memorized the entire datasets. We also observe that GPT-4 exhibits somewhat stronger evidence for memorization than GPT-3.5, which aligns with the literature that shows that larger models exhibit more memorization (Carlini et al., 2022a). Interestingly, there is no evidence of memorization of random rows on Adult and Housing .
@@
On the novel datasets, there is no evidence of memorization for GPT-3.5 and GPT-4. This is to be expected for datasets released after the model was trained. In the case of the FICO dataset, it suggests that this dataset might have been protected from inclusion in the LLM training data by the need to register prior to access. [^3] Of course, it is possible that the LLM saw the dataset during training, but there was no memorization, or our memorization tests are not powerful enough.
Table 3 shows the results of the header test and the row completion test for five different open LLMs (Groeneveld et al., 2024; Riviere et al., 2024; Dubey et al., 2024; Qwen Team, 2024). Similar to the results observed in Table 2, there is evidence for the memorization of the most popular tabular datasets in almost all LLMs. However, with the exception of the Iris dataset, there is usually only evidence for the memorization of the initial rows and not of the entire dataset.
Titanic Adult Diabetes Wine Iris Sp. Titanic FICO
OLMo 7B [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image]
Gemma2 27B [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image]
Llama3 70B [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image]
Qwen1.5 72B [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image]
Llama3.1 405B [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image] [Uncaptioned image]/ [Uncaptioned image]
Table 3: Open-source and open-weight LLMs have memorized parts of the most popular tabular datasets. The table depicts the result of the Header Test and the Row Completion Test (depicted as */*) in the simplified form used in Table 2. Detail in in Supplement D.
@@ 4 Few-Shot Learning with LLMs and Tabular Data
In the previous Section, we demonstrated that LLMs have memorized many popular tabular datasets. In this Section, we investigate the few-shot learning performance of GPT-3.5 and GPT-4 on the novel and memorized datasets. In addition, we replicate the entire analysis in this Section with Llama 3.1 70B and Gemma 2 27B. This can be found in Supplement F.
@@
Few-Shot Learning with Tabular Data System: You are a classification assistant who is an expert in tabular data, data science, and cross-sectional population surveys. […]
User: IF Age = 30, WorkClass = Private, fnlwgt = 196945, […] THEN Income =
Assistant: $>$50k [few-shot examples]
User: IF Age = 28, WorkClass = Self-emp-inc, fnlwgt = 79135, […] THEN Income =
Assistant: [Model Response]
We ask the LLM to predict the value of the target variable, given the values of the other features in the data. We select 20 few-shot examples randomly and stratify the labels of the few-shot examples to match the label occurrence in the dataset. All experiments are conducted at temperature 00. A full example prompt is given in Supplement G.
@@ 4.1 Dataset Format: Original, Perturbed, Task and Statistical
Original $0.81_{\scriptscriptstyle.01}\quad 0.96_{\scriptscriptstyle.01}$ $0.78_{\scriptscriptstyle.01}\quad 0.81_{\scriptscriptstyle.01}$ $0.74_{\scriptscriptstyle.02}\quad 0.74_{\scriptscriptstyle.02}$ $0.88_{\scriptscriptstyle.02}\quad 0.96_{\scriptscriptstyle.01}$ $0.98_{\scriptscriptstyle.01}\quad 0.99_{\scriptscriptstyle.01}$
Perturbed $0.78_{\scriptscriptstyle.01}\quad 0.82_{\scriptscriptstyle.01}$ $0.78_{\scriptscriptstyle.01}\quad 0.81_{\scriptscriptstyle.01}$ $0.73_{\scriptscriptstyle.02}\quad 0.73_{\scriptscriptstyle.02}$ $0.88_{\scriptscriptstyle.02}\quad 0.95_{\scriptscriptstyle.02}$ $0.95_{\scriptscriptstyle.02}\quad 0.95_{\scriptscriptstyle.02}$
Task $0.77_{\scriptscriptstyle.01}\quad 0.80_{\scriptscriptstyle.01}$ $0.75_{\scriptscriptstyle.01}\quad 0.79_{\scriptscriptstyle.01}$ $0.70_{\scriptscriptstyle.02}\quad 0.73_{\scriptscriptstyle.02}$ $0.87_{\scriptscriptstyle.03}\quad 0.87_{\scriptscriptstyle.03}$ $0.95_{\scriptscriptstyle.02}\quad 0.95_{\scriptscriptstyle.02}$
Statistical $0.61_{\scriptscriptstyle.02}\quad 0.65_{\scriptscriptstyle.02}$ $0.70_{\scriptscriptstyle.01}\quad 0.63_{\scriptscriptstyle.02}$ $0.68_{\scriptscriptstyle.02}\quad 0.62_{\scriptscriptstyle.02}$ $0.86_{\scriptscriptstyle.03}\quad 0.90_{\scriptscriptstyle.02}$ $0.87_{\scriptscriptstyle.03}\quad 0.92_{\scriptscriptstyle.02}$
LR / GBT 0.79   /   0.80 0.86   /   0.87 0.78   /   0.75 0.98   /   0.96 0.97   /   0.95
@@
Original $0.58_{\scriptscriptstyle.02}\quad 0.67_{\scriptscriptstyle.01}$ $0.78_{\scriptscriptstyle.01}\quad 0.78_{\scriptscriptstyle.01}$ $0.69_{\scriptscriptstyle.05}\quad 0.69_{\scriptscriptstyle.05}$ $0.58_{\scriptscriptstyle.02}\quad 0.67_{\scriptscriptstyle.01}$ $0.54_{\scriptscriptstyle.02}\quad 0.62_{\scriptscriptstyle.02}$
Perturbed $0.57_{\scriptscriptstyle.02}\quad 0.67_{\scriptscriptstyle.01}$ $0.77_{\scriptscriptstyle.01}\quad 0.78_{\scriptscriptstyle.01}$ $0.69_{\scriptscriptstyle.05}\quad 0.71_{\scriptscriptstyle.05}$ $0.58_{\scriptscriptstyle.02}\quad 0.67_{\scriptscriptstyle.02}$ $0.54_{\scriptscriptstyle.02}\quad 0.62_{\scriptscriptstyle.02}$
Task $0.59_{\scriptscriptstyle.02}\quad 0.65_{\scriptscriptstyle.02}$ $0.77_{\scriptscriptstyle.01}\quad 0.77_{\scriptscriptstyle.01}$ $0.67_{\scriptscriptstyle.05}\quad 0.71_{\scriptscriptstyle.05}$ $0.61_{\scriptscriptstyle.02}\quad 0.68_{\scriptscriptstyle.01}$ $0.54_{\scriptscriptstyle.02}\quad 0.65_{\scriptscriptstyle.01}$
Statistical $0.63_{\scriptscriptstyle.02}\quad 0.66_{\scriptscriptstyle.01}$ $0.59_{\scriptscriptstyle.02}\quad 0.57_{\scriptscriptstyle.02}$ $0.56_{\scriptscriptstyle.05}\quad 0.55_{\scriptscriptstyle.05}$ $0.60_{\scriptscriptstyle.02}\quad 0.59_{\scriptscriptstyle.02}$ $0.52_{\scriptscriptstyle.02}\quad 0.48_{\scriptscriptstyle.02}$
LR / GBT 0.78   /   0.78 0.80   /   0.80 0.76   /   0.66 0.70   /   0.69 0.64   /   0.67
@@ 4.2 Memorization leads to inflated performance estimates
Table 4 depicts the few-shot learning performance of GPT-4 and GPT-3.5 on 10 different tabular datasets. On the memorized datasets depicted in Panel A of Table 4, the performance is quite impressive. In particular, GPT-4 outperforms logistic regression on 3 out of 5 datasets when prompted with the original data. However, the predictive performance of GPT-4 and GPT-3.5 drops as we move from the original data to the perturbed data. As depicted in Figures 11(a) and 11(c), the accuracy of GPT-4 and GPT-3.5 on the memorized datasets consistently drops as we present the data in the original, perturbed, task and statistical format. In particular, GPT-4 on the task data outperforms logistic regression only on 1 out of 5 datasets, and even there, the measured difference lies within the standard error.
Panel B of Table 4 depicts the few-shot learning performance of GPT-4 and GPT-3.5 on novel datasets. In contrast to the results in Panel A, GPT-4 in the original format does not outperform logistic regression on any dataset. Moreover, we also don’t observe significant changes between the original, perturbed, and task formats. This is again depicted in Figures 11(b) and 11(d). While it is always possible that the few-shot learning performance of an LLM varies with subtleties such as the length of a feature name or the number of significant digits of a number, the results in Panel B of Table 4 indicate that GPT-4 and GPT-3.5 are fairly robust to such changes.
@@
4.3 Performance depends on feature names and variable format
4.4 Language models act as statistical predictors, but lag behind traditional algorithms
In the previous section, we have seen a significant performance drop between task and statistical, indicating that the few-shot learning performance of LLMs significantly relies on the LLM’s world knowledge (Yu et al., 2023). In this section, we study the ability of language models to act as in-context statistical learners. This relates to works using small, specifically trained language models to show that in-context learning can learn simple function classes, akin to supervised learning (Garg et al., 2022; Von Oswald et al., 2023). We investigate the question in the context of GPT-3.5 and GPT-4, similarly to Bhattamishra et al. (2024).